OpenStudy (anonymous):

Find the volume of a rectangular solid with the following dimensions: 2.3 cm x 4.6 cm x 7.8 cm

Parth (parthkohli):

Well you must multiply them.

Parth (parthkohli):

\( \color{Red}{\Rightarrow \textbf{length } \times \textbf{width} \times \textbf{height} }\)

Parth (parthkohli):

\( \color{Black}{\Rightarrow 2.3 \times 4.6 \times 7.8 }\)